• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 14:59:10 on May31,2007

    Entry number 203833

    keyword=Day Shift Summary

    Started shift finishing Q2=29fm-2.

    3934 08:34 09:11 4He 1.25M events
    3935 09:12 09:42 4He 0.64M events

    At 09:30 beam was taken away to make a pass change to a Q2 of 63fm-2

    The spectrometers have been set for the new kinematics and the trigger downloaded.